# Veri-Car: Towards Open-world Vehicle Information Retrieval

### v3

@@ Abstract
Many industrial and service sectors require tools to extract vehicle characteristics from images. This is a complex task not only by the variety of noise, and large number of classes, but also by the constant introduction of new vehicle models to the market. In this paper, we present Veri-Car, an information retrieval integrated approach designed to help with this task. It leverages supervised learning techniques to accurately identify the make, type, model, year, color, and license plate of cars. The approach also addresses the challenge of handling open-world problems, where new car models and variations frequently emerge, by employing a sophisticated combination of pre-trained models, and a hierarchical multi-similarity loss. Veri-Car demonstrates robust performance, achieving high precision and accuracy in classifying both seen and unseen data. Additionally, it integrates an ensemble license plate detection, and an OCR model to extract license plate numbers with impressive accuracy.
@@ 1 Introduction
The image retrieval problem tackled by Veri-Car is inherently open-world due to the constant introduction of new car models and modifications to existing ones. Traditional classifiers often fall short when confronted with out-of-distribution (OOD) inputs, such as car models or features that were not part of the training data. To overcome this, Veri-Car employs a combination of pre-trained models and metric learning, allowing it to generalize effectively even when encountering new and unseen data. Veri-Car labels new samples as OOD, and passes them to a human for labelling. Then, the new samples are added back into the database together with their label to allow for future classification of these examples. The retrieval and OOD detection models are general, and could be deployed in a diverse set of applications across several industries by training on datasets specific to these domains, such as retail products or surveillance datasets. In this work, we aim to provide a comprehensive pipeline for a single application. Therefore, we focus on the application of these methods to automobile detection and also include a license plate detection and recognition model, which is specific to this domain.
The approach divides the retrieval task into two main components: (1) make, model, type, and year prediction, and (2) color prediction. Although there is a lack of comprehensive datasets that include all attributes simultaneously, we leveraged the Kaggle Vehicle Color Recognition (VCoR) Dataset [1] to create an auto-labeling algorithm that merges all attributes into the same dataset. This allows us to train a single model for both components.
For the model retrieval component, Veri-Car uses the Stanford Cars 196 Dataset [2]. The methodology involves leveraging the vast knowledge of pre-trained models, like CLIP [3] and OpenCLIP [4], and fine-tuning them through a metric loss that maps similar observations closer in the embedding space. The main objective is to generate high-quality embeddings, which can be used to classify seen and unseen classes. Given that car attributes have inherent hierarchical relationships to one another, we use a hierarchical version of multi-similarity (MS) loss. This approach not only captures complex relationships among the various vehicle attributes, but also prevents the model from making drastic mistakes during inference. In addition to the retrieval of the previous properties, Veri-Car includes a color retrieval component which achieves excellent performance.
@@
Veri-Car is composed of the combination of these modules and is not an end-to-end trained system. During training, they are independent and the training of any given module does not impact the training of the others. Their unification comes only in the final stage, at the construction of the overall system. At test time, Veri-Car calls each of these modules to perform inference on a new instance independently, and returns their results in a digestible way.
The main contributions of the paper include: (1) to the best of our knowledge, the first integration of an image-based vehicle information retrieval system capable of classifying: make, type, model, year, color and license plate in an open-world setting; (2) HiMS-Min, a new hierarchical version of MS loss, that improves representations of unseen classes ;(3) a retrieval approach to color classification that improves performance over standard cross-entropy trained models; (4) a framework for training license plate recognition on purely synthetic data, that achieves high performance in the multinational domain relative to a country specific model; and (5) a system for generating synthetic data to train a license plate recognition model, either through pre-training and fine-tuning, or in a single training phase, which offers strong results in both the country specific and multinational domains, even when only a small amount of labeled country specific data is available.
@@
2 Nomenclature
  • LPD: license plate detection
  • LPR: license plate recognition
  • K-NN: K nearest neighbors
  • OCR: optical character recognition
  • OOD: out-of-distribution
  • MS: multi-similarity
@@ 3 Related Work
In this section, we will separate the related work on the three components of Veri-Car. We split the literature review as, to the extent of our knowledge, there does not exist an open source full system to which we can compare. There are several systems which tackle a related problem of vehicle extraction from natural language descriptions [9, 10, 11]. While incorporating natural language could be an extension of Veri-Car, it is not included in the present study and as such, we do not compare against such methods. There are also several proprietary systems which offer solutions to either components of the Veri-Car system or to the system as a whole. Since these solutions are not open source, they do not disclose any details of their implementation, making it impossible for us to include such works in our literature review and benchmarking. One of the main contributions of Veri-Car is to offer an open source solution in this space.
@@ 3.1 Vehicle Make and Model Recognition
This is a well studied task in the closed-set setting [12, 13, 14, 15]. However, in recent years works on OOD detection and class discovery for open-world classification of vehicles have become increasingly popular. In these settings we are not only interested in creating a simple classifier, but rather include mechanisms to detect and deal with previously unseen classes. Wolf et al. [16] leveraged a knowledge-distillation approach to include the probability of confusion between classes into the learning objective in order to improve OOD detection. Moreover, Vasquez-Santiago et al. [17] used a combination of a Gaussian Mixture Model and clustering for OOD detection and class discovery on a subset of 8 models of the VMMRdb database [18]. Furthermore, a vehicle model has inherent hierarchical information since a model is the combination of the maker, the car type (sedan, SUV, hatchback, etc.), the model name and the year it was released. Buzzelli et al. [19] posed model recognition as a hierarchical classification task by adding a car type specific classification head and a make, model and year classification head for each on car type in the dataset. Contrary to previous works, our work poses the open-world problem as a retrieval task, in which Veri-Car not only classifies seen classes correctly, but also enforces system checks to identify previously unseen classes. At the same time, the hierarchy is not used to narrow down the classification task, but to create hierarchical clusters based on common characteristics on the vehicles.
@@ 3.3 License Plate Retrieval
This is one of the most researched applications in computer vision, so it is considered, for the most part, a solved problem. License plate retrieval can be broken down into two tasks which can be performed independently or together: license plate detection (LPD) and license plate recognition (LPR). The former refers to the localization of a license plates within a larger image while the latter refers to the transcription of characters from the image of a license plate. Khan et al. [33] present a thorough review of LPR methods, including LPD methods. Here, we will focus on license plate retrieval methods which aim to provide general, rather than country specific, retrieval, but we encourage readers to examine Khan et al.’s [33] survey for a more comprehensive review of state of the art methods. They focus on neural network based methods since these tend to outperform classical methods, but note that license plate retrieval can be performed without the use of deep learning. For the same reason, we will also focus our review on deep learning based methods. Also, many of the best performing license plate retrieval methods are commercial products, and are therefore absent from the academic focused review we will discuss here.
@@ 4 Veri-Car: System Overview
In this section we describe at a high level the components of the information retrieval approach. As mentioned in Section 1, Veri-Car is composed of several independent modules. Figure 1 shows a high-level view of its architecture. During inference, a new image of a car is provided as input. First, the Car Detection module detects the bounding box of the car, and crops the image accordingly. The cropped image is then fed into the OOD Detection module, which is responsible for flagging if the image is considered in-distribution (ID) or OOD, i.e. if cars with similar attributes have been seen in the training data or not. Including the OOD detection module ensures that Veri-Car does not attempt to assign cars of unseen types to a fixed set of observed classes to which they do not belong. In the case that the image is ID, the License Plate Retrieval module and the Make, Model, Type, Year and Color Retrieval modules are responsible for extracting all necessary car attributes, namely, the license plate, make, model, type, year and color of the car. Conversely, if the image is flagged as OOD, it means that the input image is sufficiently dissimilar from the training data, and is therefore likely to belong to a new and previously unseen class of data. For example, if the system was trained in 2020 and the input is an image of a car model that came out in 2022, it would be flagged as OOD. In this case, a human-in-the-loop architecture is employed to review the image. If deemed relevant, the image is labeled and reintegrated into the Veri-Car’s database. Otherwise, if the image is irrelevant — such as in cases where it does not contain a car —, it is discarded. The next sections describe in detail these modules.
@@ 5 Make, Type, Model, Year & Color Retrieval
In this section we will explain, the architecture and training specifics for the proposed make, type, model, year and color retrieval model, as well as the dataset used. We also introduce a hierarchical version of MS loss [41], which we use to train the proposed retrieval model.
5.1 Dataset
The primary dataset used for the Make, Type, Model, Year, and Color Retrieval component is the Stanford Cars 196 Dataset [2]. This dataset provides a diverse collection of car images along with their associated labels for the make, model, vehicle type, and year. It contains 16185 images and 196 classes. Each class is fully described by the concatenation of make, model, type and year labels. We divide the samples of each class in the dataset roughly in half. Then, we choose 160 classes as a seen set of classes which we use to train the model and 36 classes as an unseen set to simulate real-life open-world conditions.
The Stanford Cars dataset does not contain a color label for the cars in the dataset. Thus, we created an automatic labeling algorithm, described in the next section, to integrate the color label into the dataset.
5.1.1 Autolabeling
We started by using the Kaggle Vehicle Color Recognition (VCoR) [1] dataset, which has 15 color categories and 10,500 images, to train a classification model with ViT-B/16 as the backbone. Next, we applied this model to the Stanford Cars dataset to predict the car color of each image. Since these predictions might be incorrect, we used the KNN+ [42] algorithm to identify and remove labels that were likely to be out-of-distribution (OOD).
After filtering out the OOD images, we combined the VCoR dataset with the correctly labeled Stanford Cars images and continued training the model. We repeated this process three times, each time using a more relaxed OOD threshold (75, 85, and 95) to gradually include more challenging samples. Finally, we selected 1,000 newly labeled samples with the lowest OOD confidence scores and manually corrected their labels if needed.
The pseudo code of the algorithm is found in Algorithm 1, where $X_{\text{labeled}}$ and $X_{\text{test}}$ refer to VCoR’s training and test set images respectively and $Y_{\text{labeled}}$ and $Y_{\text{test}}$ the corresponding labels. $X_{\text{unlabeled}}$ represents the Stanford Cars dataset images before the color label is assigned, $f_{\text{classification}}$ is a classification network, $N$ is the numbers of times the process will be repeated before stopping, and $\lambda$ is the list of progressively easier thresholds for OOD detection. We decided to make each OOD detection step more lenient as we progress with the labeling process. This approach aims to increase the distribution shift with each step, as the samples accepted with stricter OOD thresholds are those that closely resemble the data points already present in the dataset.
Table 1 shows the resulting color label distribution for the Stanford Cars dataset. We can observe that there are clear minority classes, such as pink, gold and purple, which makes it a hard dataset to solve. Especially when compared to VCoR, which has a more balanced distribution among its labels.
Algorithm 1 Autolabeling
1:$X_{\text{labeled}},X_{\text{unlabeled}},X_{\text{test}},Y_{\text{labeled}},Y_{\text{test}},N,\lambda$
2:$Y_{\text{clean}}$
3:Function $\text{TrainClassifier}(X_{1},X_{2},Y_{1},Y_{2})$: return $f$
4:Function $\text{ResumeTraining}(X_{1},X_{2},Y_{1},Y_{2},f)$: return $f$
5:Function $\text{GenLabels}(X,f)$: return $Y$
6:Function $\text{CleanLabels}(X_{1},X_{2},X_{3},Y,\lambda)$: return $Y_{2}$
7:Function $\text{AggregateLabels}(X_{1},X_{2},Y_{1},Y_{2})$ return $X_{3},Y_{3}$
8:for $n\in N$ do
9:     if $n==0$ then
10:         $f_{\text{classification}}=\text{TrainClassifier}(X_{\text{labeled}},X_{\text{test}},Y_{\text{labeled}},Y_{\text{test}})$
11:     if $n>0$ then
12:         $f_{\text{classification}}=\text{ResumeTraining}(X_{\text{labeled}},X_{\text{test}},Y_{\text{labeled}},Y_{\text{test}},f_{\text{classification}})$
13:     $Y_{\text{proposed}}=\text{GenLabels}(X_{\text{unlabeled}},f_{\text{classification}})$
14:     $Y_{\text{clean}}=\text{CleanLabels}(X_{\text{labeled}},X_{\text{test}},X_{\text{unlabeled}},Y,\lambda_{n})$
15:     $X_{\text{labeled}},Y_{\text{labeled}}=\text{AggregateLabels}(X_{\text{labeled}},X_{\text{unlabeled}},Y_{\text{labeled}},Y_{\text{clean}})$
16:return $Y_{\text{clean}}$
Table 1: Comparison of the Stanford auto-generated color labels distribution.
Color Percentage VCoR Percentage Stanford Beige 5.79 2.12 Black 5.58 16.05 Blue 10.21 7.41 Brown 7.77 3.33 Gold 2.88 0.24 Green 7.74 1.74 Grey 5.88 9.99 Orange 7.34 3.49 Pink 6.64 0.12 Purple 7.37 0.27 Red 8.76 15.48 Silver 4.98 17.03 Tan 5.50 2.32 White 5.54 18.08 Yellow 7.94 2.26
5.2 Architecture
We leverage pre-trained models, specifically OpenCLIP’s ViT-B/16 [43] trained on LAION-2B [44], as the backbone to our model architecture. CLIP [3] models are known for their ability to create compressed representations (embeddings) of images that capture intricate details and relationships. CLIP’s learning objective maximizes the similarity between a vector representation of an image, and its corresponding text description in the embedding space. The combination of its training framework, and the size of the dataset make it a great initialization point.
In our experiments we use the image encoder and discard the text encoder provided in CLIP [3]. Since adding color to the same embedding space could result in the network associating car colors with specific models/manufacturers, we place two separate linear layers of size 128 on top of the image encoder to learn the lower dimensional embeddings to be used for $K$-NN classification (Figure 2). This architecture allows us to learn high-quality embedding representations of images with limited samples-per-class while minimizing correlations between colors and models.
Figure 2: Make, Type, Model, Year and Color Retrieval model.
@@ 5.3 Multi-Similarity Loss
The core of the model’s retrieval process is based on metric learning, which aims to map similar observations closer in the embedding space. The MS loss was introduced by Wang et al. [41] to enhance the performance of deep metric learning by effectively leveraging both the hardest-positive and hardest-negative pairs within each batch. This allows us to choose a positive and negative pair for each image in the batch with the most relevant information, optimizing the model to distinguish between similar and dissimilar samples effectively. The negative pair for an anchor image $i$ is defined as the image $k$, from a different class, with a higher similarity $s_{ik}$ than the hardest positive (image from the same class with lowest similarity). The positive pair is defined as the image $j$, belonging to the same class as $i$, with a lower similarity $s_{ij}$ than the hardest negative (image from a different class with the highest similarity). In the context of model retrieval, positive pairs consist of pairs of observations which belong to the same car make-type-model-year class, and negative pairs belong to two different car make-type-model-year classes. The formula for MS loss is shown in Equation 1, and ensures that the model learns to create clusters for samples of cars of the same characteristics, while maintaining a clear separation from samples of cars of different characteristics.
@@ 5.4 Hierarchical Multi-Similarity Loss
Since the nature of our data is hierarchical, we created a hierarchical version of MS loss [41]. A hierarchical MS loss identifies positive and negative pairs at all hierarchical levels for all samples in the batch aiming to create super-clusters of semantically similar samples. Formally, given an anchor image $i$, a set of hierarchy levels $L$ and a set of positive and negative pairs $Z_{l}(i)\,\forall\,l\in L$, where $Z_{l}(i)=P_{l}(i)\cup N_{l}(i)$, positive pairs $P_{l}(i)$ will be pushed closer the further down the hierarchy we go while negative pairs $N_{l}(i)$ will be pushed farther away as we move up in the hierarchy. Figure 3 shows a visual representation of the hierarchical label structure created for the Make, Type, Model, Year component of the model. Samples at the leaves of the tree have the same make, type, model and year. Color was excluded from this hierarchy. This approach ensures that the model accurately learns the relationships among different vehicle attributes.
@@
The hierarchical MS loss was created following Zhang et al.’s  [45] view on hierarchical contrastive learning where they use the inherent hierarchical tree-like relationships among the labels to optimize the model. Their proposed loss is shown in Equation 2 and ensures that as we traverse the label tree downwards the loss at lower levels is never lower than at higher levels. The loss function is shown in Equation 2.
@@
The above formulation forces lower levels of the hierarchy to have at least the same importance as that of higher levels. However, it might push sub-clusters too close to one another. Thus, we flipped the hierarchy, and forced higher levels of the hierarchy to have, at most, the same importance as lower levels. Such a formulation should create looser clusters as we go up the hierarchy, but still tight clusters at the leaves at no extra computational cost. Thus, the new loss function is defined in Equation 3.
$\mathcal{L}_{HiMS-Min}=\sum_{L=l,...,1}\frac{1}{|L|}\sum_{i=1}^{B}\frac{\lambda_{l}}{|Z(i)|}\sum_{z_{l}\in\mathcal{Z}(i)_{l}}\text{min}(\mathcal{L}_{MS}(i,z^{i}_{l}),\mathcal{L}^{min}_{MS}(l+1))$ (3)
where $L_{MS}$ is the MS loss in Equation 1 and $\lambda=\text{exp}(1/l)$.
@@
5.5 Model Configurations

### v4

@@ Abstract
Many industrial and service sectors require tools to extract vehicle characteristics from images. This is a complex task not only by the variety of noise, and large number of classes, but also by the constant introduction of new vehicle models to the market. In this paper, we present Veri-Car, an information retrieval integrated approach designed to help on this task. It leverages supervised learning techniques to accurately identify the make, type, model, year, color, and license plate of cars. The approach also addresses the challenge of handling open-world problems, where new car models and variations frequently emerge, by employing a sophisticated combination of pre-trained models, and a hierarchical multi-similarity loss. Veri-Car demonstrates robust performance, achieving high precision and accuracy in classifying both seen and unseen data. Additionally, it integrates an ensemble license plate detection, and an OCR model to extract license plate numbers with impressive accuracy.
@@ 1 Introduction
The image retrieval problem tackled by Veri-Car is inherently open-world due to the constant introduction of new car models and modifications to existing ones. Traditional classifiers often fall short when confronted with out-of-distribution (OOD) inputs, such as car models or features that were not part of the training data. To overcome this, Veri-Car employs a combination of pre-trained models and metric learning, allowing it to generalize effectively even when encountering new and unseen data. Veri-Car labels new samples as OOD, and passes them to a human for labelling. Then, the new samples are added back into the database together with their label to allow for future classification of these examples. The retrieval and OOD detection models are general, and could be deployed in a diverse set of applications across several industries by training on datasets specific to these domains, such as retail products datasets.
The approach divides the retrieval task into two main components: (1) make, model, type, and year prediction, and (2) color prediction. Although both components could be merged into one, they were separated due to a lack of comprehensive datasets that include all attributes simultaneously.
For the model retrieval component, Veri-Car uses the Stanford Cars 196 Dataset [1]. The methodology involves leveraging the vast knowledge of pre-trained models, like CLIP [2] and OpenCLIP [3], and fine-tuning them through a metric loss that maps similar observations closer in the embedding space. The main objective is to generate high-quality embeddings, which can be used to classify seen and unseen classes. Given that car attributes have inherent hierarchical relationships to one another, we use a hierarchical version of multi-similarity loss. This approach not only captures complex relationships among the various vehicle attributes, but also prevents the model from making drastic mistakes during inference.
In addition to the retrieval of the previous properties, Veri-Car includes a color retrieval component trained using the Kaggle Vehicle Color Recognition (VCoR) Dataset [4]. Employing a similar architecture and multi-similarity loss, this model achieves excellent performance.
@@
The main contributions of the paper include: (1) to the best of our knowledge, the first integration of an image-based vehicle information retrieval system capable of classifying: make, type, model, year, color and license plate in an open-world setting; (2) HiMS-Min, a new hierarchical version of Multi-Similarity loss, that improves representations of unseen classes ;(3) a retrieval approach to color classification that improves performance over standard cross-entropy trained models; (4) a framework for training license plate recognition on purely synthetic data, that achieves high performance in the multinational domain relative to a country specific model; and (5) a system for generating synthetic data to train a license plate recognition model, either through pre-training and fine-tuning, or in a single training phase, which offers strong results in both the country specific and multinational domains, even when only a small amount of labeled country specific data is available.
@@ 2 Related Work
In this section, we will separate the related work on the three components of Veri-Car.
@@ 2.1 Vehicle Make and Model Recognition
This is a well studied task in the closed-set setting [9, 10, 11, 12]. However, in recent years works on out-of-distribution detection and class discovery for open-world classification of vehicles have become increasingly popular. In these settings we are not only interested in creating a simple classifier, but rather include mechanisms to detect and deal with previously unseen classes. Wolf et al. [13] leveraged a knowledge-distillation approach to include the probability of confusion between classes into the learning objective in order to improve OOD detection. Moreover, Vasquez-Santiago et al. [14] used a combination of a Gaussian Mixture Model and clustering for OOD detection and class discovery on a subset of 8 models of the VMMRdb database [15]. Furthermore, a vehicle model has inherent hierarchical information since a model is the combination of the maker, the car type (sedan, SUV, hatchback, etc.), the model name and the year it was released. Buzzelli et al. [16] posed model recognition as a hierarchical classification task by adding a car type specific classification head and a make, model and year classification head for each on car type in the dataset. Contrary to previous works, our work poses the open-world problem as a retrieval task, in which Veri-Car not only classifies seen classes correctly, but also enforces system checks to identify previously unseen classes. At the same time, the hierarchy is not used to narrow down the classification task, but to create hierarchical clusters based on common characteristics on the vehicles.
@@ 2.3 License Plate Retrieval
This is one of the most researched applications in computer vision, so it is considered, for the most part, a solved problem. Licence plate retrieval can be broken down into two tasks which can be performed independently or together: license plate detection (LPD) and license plate recognition (LPR). The former refers to the localization of a license plates within a larger image while the latter refers to the transcription of characters from the image of a license plate. Khan et al. [30] present a thorough review of LPR methods, including LPD methods. Here, we will focus on license plate retrieval methods which aim to provide general, rather than country specific, retrieval, but we encourage readers to examine Khan et al.’s [30] survey for a more comprehensive review of state of the art methods. They focus on neural network based methods since these tend to outperform classical methods, but note that license plate retrieval can be performed without the use of deep learning. For the same reason, we will also focus our review on deep learning based methods. Also, many of the best performing license plate retrieval methods are commercial products, and are therefore absent from the academic focused review we will discuss here.
@@ 3 Veri-Car: System Overview
In this section we describe at a high level the components of the information retrieval approach. As mentioned in Section 1, Veri-Car is composed of several independent modules. Figure 1 shows a high-level view of its architecture. During inference, a new car image is provided as input. First, the Car Detection module detects the bounding box of the car, and crops the image accordingly. The cropped image is then fed into the OOD Detection module, which is responsible for flagging if the image is considered in-distribution or out-of-distribution, i.e. if cars of similar attributes have been seen in the training data or not. In the case that the image is in-distribution, the License Plate Retrieval, Make, Model, Type, Year Retrieval and the Color Retrieval modules are responsible for extracting all necessary car attributes, namely, the license plate, make, model, type, year and color of the car. Conversely, if the image is flagged as out-of-distribution, a human-in-the-loop architecture is employed to review the image. If deemed relevant, the image is labeled and reintegrated into the Veri-Car’s database. Otherwise, if the image is irrelevant — such as in cases where it does not contain a car —, it is discarded. The next sections describe in detail these modules.
@@ 4 Make, Type, Model, Year & Color Retrieval
In this section, we integrate both the Make, Type, Model, Year Retrieval and the Color Retrieval models. We discuss these two models together because of their similarities. We use $k$-NN as the main classification algorithm on the extracted embeddings of vehicle images as explained below in the Architecture subsection. At the same time, we introduce a hierarchical version of Multi-similarity loss [38] used to train the Make, Type, Model, Year Retrieval model while standard Multi-similarity loss [38] is used to train the Color Retrieval model. The only reason for keeping these two models distinct was the lack of comprehensive datasets combining all necessary labels.
@@ 4.1 Multi-Similarity Loss
The core of the model’s retrieval process is based on metric learning, which aims to map similar observations closer in the embedding space. The multi-similarity loss was introduced by Wang et al. [38] to enhance the performance of deep metric learning by effectively leveraging both the hardest-positive and hardest-negative pairs within each batch. This allows us to choose a positive and negative pair for each image in the batch with the most relevant information, optimizing the model to distinguish between similar and dissimilar samples effectively. The negative pair for an anchor image $i$ is defined as the image $k$, from a different class, with a higher similarity $s_{ik}$ than the hardest positive (image from the same class with lowest similarity). The positive pair is defined as the image $j$, belonging to the same class as $i$, with a lower similarity $s_{ij}$ than the hardest negative (image from a different class with the highest similarity). In the context of model retrieval, positive pairs consist of pairs of observations which belong to the same car make-type-model-year class, and negative pairs belong to two different car make-type-model-year classes. The formula for multi-similarity loss is shown in Equation 1, and ensures that the model learns to create clusters for samples of cars of the same characteristics, while maintaining a clear separation from samples of cars of different characteristics.
@@ 4.2 Hierarchical Multi-Similarity Loss
Since the nature of our data is hierarchical, we created a hierarchical version of multi-similarity loss [38]. A hierarchical multi-similarity loss identifies positive and negative pairs at all hierarchical levels for all samples in the batch aiming to create super-clusters of semantically similar samples. Formally, given an anchor image $i$, a set of hierarchy levels $L$ and a set of positive and negative pairs $Z_{l}(i)\,\forall\,l\in L$, where $Z_{l}(i)=P_{l}(i)\cup N_{l}(i)$, positive pairs $P_{l}(i)$ will be pushed closer the further down the hierarchy we go while negative pairs $N_{l}(i)$ will be pushed farther away as we move up in the hierarchy. Figure 2 shows a visual representation of the hierarchical label structure created for the Make, Type, Model, Year Retrieval model. Samples at the leaves of the tree have the same make, type, model and year. This approach ensures that the model accurately learns the relationships among different vehicle attributes.
@@
The hierarchical multi-similarity loss was created following Zhang et al.’s  [39] view on hierarchical contrastive learning where they use the inherent hierarchical tree-like relationships among the labels to optimize the model. Their proposed loss is shown in Equation 2 and ensures that as we traverse the label tree downwards the loss at lower levels is never lower than at higher levels. The loss function is shown in Equation 2.
@@
The above formulation forces lower levels of the hierarchy to have at least the same importance as that of higher levels. However, it might push sub-clusters too close to one another. Thus, we flipped the hierarchy, and forced higher levels of the hierarchy to have, at most, the same importance as lower levels. Such a formulation should create looser clusters as we go up the hierarchy, but still tight clusters at the leaves. Thus, the new loss function is defined in Equation 3.
$\mathcal{L}_{HiMS-Min}=\sum_{L=l,...,1}\frac{1}{|L|}\sum_{i=1}^{B}\frac{\lambda_{l}}{|Z(i)|}\sum_{z_{l}\in\mathcal{Z}(i)_{l}}\text{min}(\mathcal{L}_{MS}(i,z^{i}_{l}),\mathcal{L}^{min}_{MS}(l+1)$ (3)
where $L_{MS}$ is the Multi-similarity loss in Equation 1 and $\lambda=\text{exp}(1/l)$.
@@
4.3 Make, Type, Model, Year Retrieval model
4.3.1 Dataset
The primary dataset used for the Make, Type, Model, Year Retrieval component is the Stanford Cars 196 Dataset [1]. This dataset provides a diverse collection of car images along with their associated labels for the make, model, vehicle type, and year. The dataset is composed of 16185 images and 196 classes. Each class is fully described by the concatenation of make, model, type and year labels. We divide the samples of each class in the dataset roughly in half. Then, we choose 160 classes as a seen set of classes which we use to train the model and 36 classes as an unseen set to simulate real-life open-world conditions.
4.3.2 Architecture
We leverage pre-trained models, specifically OpenCLIP’s ViT-B/16 [41] trained on LAION-2B [42] as the backbone to our model architecture. CLIP [2] models are known for their ability to create compressed representations (embeddings) of images that capture intricate details and relationships. CLIP’s learning objective maximizes the similarity between a vector representation of an image, and its corresponding text description in the embedding space. The combination of its training framework, and the size of the dataset make it a great initialization point.
In our experiments we use the image encoder and discard the text encoder provided in CLIP [2]. Then, a linear layer of size 128 is placed on top of the image encoder to learn a lower dimensional embedding to be used for $k$-NN classification. This architecture allows us to learn high-quality embedding representations of images with limited samples-per-class.
4.3.3 Model Configurations